OpenStudy (anonymous):

Which is closest to the 22nd term in the following sequence? 16.4, –14.76, 13.284, –11.9556

OpenStudy (anonymous):

16.4

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

i gave u answer i have the florida virtual school i had the same question 2day

OpenStudy (anonymous):

i passed the test and git this question right

OpenStudy (anonymous):

1.62?

OpenStudy (whpalmer4):

What do you get if you do the following divisions? 16.4/-14.76= -14.76/13.284= 13.284/-11.9556= Remember that a geometric sequence is one where each term produces the next term by multiplication by a constant. Nothing says the constant has to be positive!

OpenStudy (whpalmer4):

If you have a sequence \(\{a_k\}, k = 1, 2, ... \text{and } a_k = a_0 r^k\). \[a_0 = 16.4\]\[r = (-14.76/16.4)\]\[a_{22} = a_0 r^{22} = \] Looks like the ratio is supposed to be -9/10.